# RUMOR: Reinforcement learning for Understanding a Model of the Real World for Navigation in Dynamic Environments

### v1

@@ 4 Reinforcement learning setup
where $\gamma$ is a discount factor to balance the present value of future rewards and $V(\mathbf{s}_{t})$ is the value function, that represents the expected cumulative reward starting from state $\mathbf{s}_{t}$ and following the policy $\pi(\mathbf{s}_{t})$:
@@
4.1 State space
The state $\mathbf{s}_{t}\in\mathcal{S}$ is the unobservable true condition of the system that drives the dynamics of the environment. In our problem, we consider that there is a robot with radius $r_{robot}$ position $\mathbf{p}_{t}$ and velocity $\mathbf{u}_{t}$ at time $t$ that drives towards a goal $\mathbf{g}$. The rest of the state is the set of points occupied by the set of static and dynamic obstacles $\mathscr{O}_{\forall,t}$ and the radius $\mathbf{r}_{obs}\in\mathbb{R}^{M}$ and velocity $\mathscr{V}_{t}=\{(v_{i},\omega_{i})\forall i\in[0\dots M]\}$ of the $M$ obstacles that are dynamic:
@@ 4.1 State space
4.2 Action space
The action $\mathbf{a}_{t}\in\mathcal{A}$ should account for the kinematics and dynamics of the robot stated in Equation 3 and Equation 4. To do so, we use a continuous action set of two values $\mathbf{a}_{t}\in[0,1]^{2}=(a_{1,t},a_{2,t})$ and relate it to the constraints. We first define the four corners of the dynamic window of Equation 4:
@@ 4.2 Action space
Then, define two vectors to form a basis in the velocity space:
@@
We can compute a velocity that accounts for the acceleration constraints as:
To account for the dynamic restriction of Equation 3, $\mathbf{a}_{t}$ maximum value is bounded: $a_{1,t}\in[0,a_{1,t}^{max}]$ and $a_{2,t}\in[0,a_{2,t}^{max}]$. It will be bounded if the dynamic window intersects with $l_{1}$ or $l_{2}$. In the velocity space, the lines that pass through $\mathbf{u}_{t}^{down}=(\omega_{t}^{down},v_{t}^{down})$ and have the direction of $\mathbf{b}_{1,t}$ and $\mathbf{b}_{2,t}$ are $l_{1}^{\prime}$ and $l_{2}^{\prime}$, respectively:
@@
And the intersection points between $l_{1}$ and $l_{1}^{\prime}$, and $l_{2}$ and $l_{2}^{\prime}$ are $P_{1}(\omega_{1,p},v_{2,p})$ and $P_{2}(\omega_{2,p},v_{2,p})$, with:
$\begin{split}&\omega_{1,p}=\frac{1}{2}(\frac{\omega_{max}v_{t}^{down}}{v_{max}}+\omega_{t}^{down}-\omega_{max})\\ &v_{1,p}=\frac{1}{2}(\frac{v_{max}\omega_{t}^{down}}{\omega_{max}}+v_{t}^{down}-v_{max})+v_{max}\\ &\omega_{2,p}=\frac{1}{2}(-\frac{\omega_{max}v_{t}^{down}}{v_{max}}+\omega_{t}^{down}+\omega_{max})\\ &v_{2,p}=\frac{1}{2}(-\frac{v_{max}\omega_{t}^{down}}{\omega_{max}}+v_{t}^{down}-v_{max})+v_{max}\\ \end{split}$ (17)
The action $\mathbf{a}_{t}$ will be bounded only if the distance between $\mathbf{u}_{t}^{down}$ and $P_{1}$ and $P_{2}$ is smaller than $\mathbf{b}_{1,t}$ or $\mathbf{b}_{2,t}$, since the dynamic window will be in collision with $l_{1}$ or $l_{2}$. Therefore:
@@
The elements used in the velocity computation are represented in Figure 4. It may be seen how the result of the linear combination of Equation 15, with the action bounded by Equation LABEL:eq:action-bounds, produces a velocity inside the rhombus defined by Equation 4 and respects Equation 3.
Figure 4: Graphical representation of the notation used in the action space configuration. $\mathbf{u}_{t}$, $\mathbf{u}_{t}^{up}$, $\mathbf{u}_{t}^{down}$, $\mathbf{u}_{t}^{left}$ and $\mathbf{u}_{t}^{right}$ are represented with gray points, $l_{1}$ and $l_{2}$ with black lines, $l_{1}^{\prime}$ and $l_{2}^{\prime}$ with dotted lines, $\mathbf{b}_{1,t}$ and $\mathbf{b}_{2,t}$ with blue arrows and $P_{2}$ with a red point.
4.3 Observation space, observation model and transition model
@@
The robot is modeled to have a 2-D LiDAR laser sensor. It estimates $\hat{\mathbf{p}}_{t}$ and $\hat{\mathbf{u}}_{t}$ with the laser data and the odometry, using Adaptive Monte Carlo Localization (AMCL) [38]. An extended version of the work proposed by [39] is used to detect static and dynamic obstacles and estimate $\hat{\mathscr{O}}_{\forall,t}$, $\hat{\mathbf{r}}_{obs}$ and $\hat{\mathscr{V}}_{t}$ from the 2-D LiDAR measurements. It detects circular and linear obstacles from the scans, associating obstacles of any shape to circles or a set of segments, and tracks and estimates the velocity of the circular obstacles using an algorithm based on Kalman Filter. This kind of conditions makes the robot face simulated occlusions and estimation errors. In the simulations, we only consider dynamic circular obstacles to be able to compare out method with other approaches that assume that all obstacles are circular, but our approach could handle obstacles with any shape.
@@ 4.4 Reward function
where $d_{t,goal}=||\mathbf{g}-\mathbf{u}_{t}||$ and $\Delta{d_{t,goal}}=d_{t,goal}-d_{t-1,goal}$. The robot receives a reward of $r_{goal}=15$, when it reaches the goal within a certain threshold; and a negative reward $r_{collision}=-15$ when it collides. Reward shaping is used to accelerate training in non-terminal states by encouraging the agent to get closer to the goal ($-r_{dist}\Delta{d_{t,goal}}$ with $r_{dist}=2.5$), and by penalizing the agent if it is too close to an obstacle with:
@@
5 Network
We propose using a Soft Actor Critic (SAC) [19] algorithm to solve the problem, and use the DOVS model as part of the input of the network to extract the dynamism of the environment. The implementation of Stable Baselines 3 [42] is used for the SAC. SAC is an off-policy DRL algorithm that uses an actor to determine the policy and a to estimate critic the state-action values, and introduces entropy maximization to help exploration; it is known for its stability and sample efficiency.
@@ 5 Network
$\mathbf{o}_{t,state}=\left\{\hat{\mathbf{u}}_{t},\hat{d}_{t,goal},\hat{\theta}_{t,goal},\hat{\mathscr{O}}_{t,dist},\hat{\mathscr{O}}_{t,\theta},\hat{\mathscr{O}}_{v,t},\hat{\mathscr{O}}_{t,dir}\right\},$ (23)
where $\hat{d}_{t,goal}$ and $\hat{\theta}_{t,goal}$ are the distance and angle to the goal with respect to the estimated heading angle of the robot ($\theta_{t}$), $\hat{\mathscr{O}}_{t,dist}$ the estimated distance of the robot to the closest obstacle, $\hat{\mathscr{O}}_{t,\theta}$ the estimated angle to it, $\hat{\mathscr{O}}_{v,t}$ its estimated velocity and $\hat{\mathscr{O}}_{t,dir}$ its estimated heading angle. Even though the information of these four last variables is already encoded in DOVS, they are included to give more information to the robot in case there is an imminent collision.
@@
The output of both streams is concatenated and fed to an LSTM layer [23], $\psi_{LSTM}$, to save information of previous observation and keep temporal dependencies. The future of the environment is already considered by DOVS, but considering the past could help the network understand changes in the obstacles behavior and make it more robust to estimation errors or occlusions. Finally, a fully connected layer with ReLU activation function, $\psi_{FC}$, to produce the final encoded observation $\mathbf{o}_{t,enc}$:
@@ 6 Experiments
6.1 Training details
The agent has been trained in a modified version of the Stage simulator [43], in a scenario with dynamic obstacles the robot has to avoid to reach the goal. This simulator is conveniently integrated in ROS to use AMCL for localization and realistically simulates a LiDAR sensor. Random positions are sampled for the robot, its goal and the other obstacles. The obstacles tried to avoid each other using the ORCA algorithm, but they are not able to see the robot to make it perform the whole obstacle avoidance maneuvers, as stated in Section 4.3. The obstacles have a predefined random linear and angular velocity that try to follow, between 0.14 and 0.71 m/s. The robot faces scenarios with increasing number of dynamic obstacles, up to 14, and increasing starting distance to the goal during a first training stage of 1000 episodes. In this way, it learns to reach the goal progressively by using curriculum learning, from simple to more complex tasks. After that stage, it faces 9000 scenarios with a random number of obstacles up to 14. The minimum distance between the robot and the goal is set to 6 m. The robot velocity limits are set to $v_{max}=0.7$ m/s and $\omega_{max}=\pi$ rad/s and the maximum acceleration to $a_{max}=0.3$ m/s², similar to the ones of a Turtlebot 2 platform.
@@ 6.2 Simulation experiments
In this part, we conducted a series of quantitative experiments to evaluate the performance of our proposed model. We compared our method, RUMOR, with ORCA [11], S-DOVS [1], SARL [26] and RE3-RL [29], in order to test other model-based and DRL-based planners of the state of the art for navigation in dynamic environments. Nevertheless, the planners consider different constraints, as stated in Table 1. Only our planner and S-DOVS consider the constraints stated in Equation 3 and Equation 4. We propose an alternative version of our planner that does not respect those constraints for a more fair comparison, called NR-RUMOR (no-restrictions). It uses as the action space velocities ranging from the minimum and maximum limits, similar to SARL and RE3-RL.
We tested the planners in the same 100 random scenarios with low and high occupancy, with 6 and 12 dynamic obstacles, respectively. The 85% of the obstacles are set to be dynamic and the same set of scenarios is used for every version. We also tested two different perception settings: using the ground truth position of the dynamic obstacles given by the simulator (absolute perception) and using the Kalman Filter-based obstacle detector (obstacle tracker) previously referenced in Section 4.3, thus facing occlusions and estimation errors. The DRL policies (both ours and those of the state of the art) have been trained in the same settings as in the testing experiments, i.e., there are two versions of each of the DRL planners, one trained using absolute perception and the other using the obstacle tracker. The results obtained are plotted in Figure 7, where the success rates, navigation times path lenghts and mean linear velocities recorded are plotted.
Motion DRL Differential Kinodynamics
Planner drive
NR-RUMOR (ours)
RUMOR (ours)
S-DOVS
ORCA
SARL
RE3-RL
Table 1: Differences between motion planners, where DRL refers to being a DRL-based planner, differential drive if they outup differential drive commands and kinodynamics if they consider acceleration constraints or other dynamic and kinematic restrictions.
Absolute perception (6) Obstacle tracker (6)
@@
(a1) Success rate (a2) Success rate
@@
(b1) Navigation time (b2) Navigation time
@@
(c1) Path length (c2) Path length
@@
(d1) Mean v (d2) Mean v
Absolute perception (12) Obstacle tracker (12)
@@
(a3) Success rate (a4) Success rate
@@
(b3) Navigation time (b4) Navigation time
Refer to caption Refer to caption
(c3) Path length (c4) Path length
Refer to caption Refer to caption
(d3) Mean v (d4) Mean v
@@
Figure 7: Navigation metrics for dynamic scenarios of 6 (left) and 12 (right) obstacles using absolute perception and the obstacle tracker.
We mainly evaluate the planners using the two navigation metrics directly encoded in the reward function, which are the success rate (number of times the robot have reached the goal without collisions or timeouts) and the navigation time. There is a trade-off between both metrics. Trying to achieve low navigation time leads to risky collision avoidance maneuvers, which leads to lower success rates. In addition, the task is complex, and there are scenarios where the randomness in the positions and velocities of the obstacles makes it impossible to avoid collisions. For example, specially in dense scenarios, the robot may face trapping situations where several obstacles run towards it.
@@
The success rates show, as previously stated by other works, that both model-based planners performance is worse than DRL-based ones. ORCA exhibits a lower navigation time than the other planners, probably due to the fact that it is a holonomic planner and does not respect differential-drive motion, but the success rate of both ORCA and S-DOVS are clearly the lowest in all settings. These planners model the environment in every time step, but they are not robust when errors or deep changes in the surroundings occur.
@@
Regarding the perception settings, every planner shows higher success rates and lower navigation times with absolute perception, due to the absence of partial observability and estimation errors. It is way more complex to predict the future of the environment when facing perception problems. Therefore, their relative performance within the same settings should be compared. It is interesting to see that there is not consistency in the performance of SARL and RE3-RL when changing the perception settings. When comparing them using absolute perception, SARL outperforms RE3-RL in both success rate and navigation time, while when using the obstacle tracker is the other way around. This is probably due to the difference in the complexity of their networks and the exploration process. Using intrinsic rewards to explore, as done in RE3-RL, may degrade the performance of the planner in the presence of erroneous observations. In contrast, both NR-RUMOR and RUMOR show a very consistent relative performance no matter the perception settings.
In the low occupancy experiments of Figure 7 (1-2), the performance of NR-RUMOR is similar to the best DRL planner in every perception settings. These scenarios are relatively easy and standard end-to-end approaches show good performance. NR-RUMOR slightly outperforms RE3-RL in both success and time with absolute perception, while with the obstacle tracker, SARL is the method which slightly outperforms NR-RUMOR. Nevertheless, it is in the high occupancy experiments of Figure 7 (3-4) where the superior performance of NR-RUMOR is stated. It clearly surpasses RE3-RL in terms of success and time with absolute perception, while, with the obstacle tracker, the navigation time regarding SARL is comparable, but the success rate is not. Two key components of the system explains this fact. First, the input of RUMOR network is the DOVS model of the environment. With 12 surrounding dynamic obstacles, there are unlimited number of possible scenario configurations, and the agent can not be trained to have a robust performance in all of them. On the contrary, DOVS represents the environment in similar terms no matter the scenario, leading to a deeper understanding of the velocity space. When the robot learns to interpret DOVS, it knows how to avoid collisions regardless the amount, positions or velocities of the obstacles. Second, the memory layer of RUMOR network allows it to keep track of previous observations. This is important to overcome partial observability challenges and those derived from sudden changes in the obstacles behavior. Overall, considering every combination in the perception and occupancy settings, NR-RUMOR consistently shows the best performance when compared with the other unconstrained or model-based methods.

### v2

@@ 4.2 Reinforcement learning setup
where $\gamma$ is a discount factor to balance the present value of future rewards, and $V(\mathbf{s}_{t})$ is the value function, which represents the expected cumulative reward starting from state $\mathbf{s}_{t}$ and following the policy $\pi(\mathbf{s}_{t})$:
@@
4.2.1 State space
The state $\mathbf{s}_{t}\in\mathcal{S}$ is the unobservable true condition of the system that drives the dynamics of the environment. In our problem, we consider that there is a robot with radius $r_{robot}$ position $\mathbf{p}_{t}$ and velocity $\mathbf{u}_{t}$ at time $t$ that drives towards a goal $\mathbf{g}$. The rest of the state is the set of points occupied by the set of static and dynamic obstacles $\mathscr{O}_{\forall,t}$, the radius $\mathbf{r}_{obs}\in\mathbb{R}^{M}$ and velocity $\mathscr{V}_{t}=\{(v_{i},\omega_{i})\forall i\in[0\dots M]\}$ of the $M$ obstacles that are dynamic:
@@ 4.2.1 State space
4.2.2 Action space
The action $\mathbf{a}_{t}\in\mathcal{A}$ should account for the kinematics and dynamics of the robot stated in Equation 3 and Equation 4. To do so, we geometrically design an action space that ensures the constraints are not violated by construction. This construction is represented in Figure 4. Having that the robot has the velocity $\boldsymbol{u}_{t}$ in the current control period, $\boldsymbol{u}_{t+1}$ must not be outside the green rhombus (the dynamic window of Equation 4) and must not surpass $l_{1}$ nor $l_{2}$ (Equation 3). The idea of the action space is solving this problem by selecting velocities that respect the restrictions by definition. First, we construct an action space that make velocities be always in the dynamic window. Second, we constraint the dynamic window if it intersects with constraints in Equation 3.
We define a continuous action space of two values $\mathbf{a}_{t}=(a_{1,t},a_{2,t})\in[0,1]^{2}$ that linearly combine $\mathbf{b}_{1,t}$ and $\mathbf{b}_{2,t}$, to compute velocities that always remain inside the dynamic window:
Figure 4: Graphical representation of the notation used in the action space configuration, attaining for restrictions previously represented in Figure 2. $\mathbf{u}_{t}$, $\mathbf{u}_{t}^{up}$, $\mathbf{u}_{t}^{down}$, $\mathbf{u}_{t}^{left}$ and $\mathbf{u}_{t}^{right}$ are represented with gray points, $l_{1}$ and $l_{2}$ with black lines, $l_{1}^{\prime}$ and $l_{2}^{\prime}$ with dotted lines, $\mathbf{b}_{1,t}$ and $\mathbf{b}_{2,t}$ with blue arrows, and $Q_{1}$ and $Q_{2}$ with red points.
To do so, the four corners of the dynamic window are calculated:
@@ 4.2.2 Action space
Using them, the two vectors to form a basis in the velocity space are defined:
@@
The kynematic restriction of Equation 3 ($l_{1}$ and $l_{2}$ in Figure 4) are considered by limiting $\boldsymbol{a}_{t}$ maximum value: $a_{1,t}\in[0,a_{1,t}^{max}]$ and $a_{2,t}\in[0,a_{2,t}^{max}]$. This limit is only needed if the dynamic window intersects with $l_{1}$ or $l_{2}$. In the velocity space, the lines that pass through $\mathbf{u}_{t}^{down}=(\omega_{t}^{down},v_{t}^{down})$ and have the direction of $\mathbf{b}_{1,t}$ and $\mathbf{b}_{2,t}$ are $l_{1}^{\prime}$ and $l_{2}^{\prime}$, respectively:
@@
And the intersection points between $l_{1}$ and $l_{1}^{\prime}$, and $l_{2}$ and $l_{2}^{\prime}$ are $Q_{1}(\omega_{1,q},v_{1,q})$ and $Q_{2}(\omega_{2,q},v_{2,q})$, with:
$\begin{split}&\omega_{1,q}=\frac{1}{2}(\frac{\omega_{max}v_{t}^{down}}{v_{max}}+\omega_{t}^{down}-\omega_{max})\\ &v_{1,q}=\frac{1}{2}(\frac{v_{max}\omega_{t}^{down}}{\omega_{max}}+v_{t}^{down}-v_{max})+v_{max}\\ &\omega_{2,q}=\frac{1}{2}(-\frac{\omega_{max}v_{t}^{down}}{v_{max}}+\omega_{t}^{down}+\omega_{max})\\ &v_{2,q}=\frac{1}{2}(-\frac{v_{max}\omega_{t}^{down}}{\omega_{max}}+v_{t}^{down}-v_{max})+v_{max}\\ \end{split}$ (17)
The action $\mathbf{a}_{t}$ will be bounded only if the distance between $\mathbf{u}_{t}^{down}$ and $Q_{1}$ and $Q_{2}$ is smaller than $\mathbf{b}_{1,t}$ or $\mathbf{b}_{2,t}$, since the dynamic window will be in collision with $l_{1}$ or $l_{2}$. Therefore:
@@
4.2.3 Observation space, observation model and transition model
@@
The robot is modeled to have a 2-D LiDAR laser sensor. It estimates $\hat{\mathbf{p}}_{t}$ and $\hat{\mathbf{u}}_{t}$ with the laser data and the odometry, using Adaptive Monte Carlo Localization (AMCL) [56]. An extended version of the work proposed by [57] is used to detect static and dynamic obstacles and estimate $\hat{\mathscr{O}}_{\forall,t}$, $\hat{\mathbf{r}}_{obs}$ and $\hat{\mathscr{V}}_{t}$ from the 2-D LiDAR measurements. It detects circular and linear obstacles from the scans, by using a grouping and splitting algorithm, associating obstacles of any shape to circles or a set of segments. The circular obstacles are tracked in time with a Kalman filter, wich also estimates its velocity assuming constant velocity in the update step of the Kalman filter. We refer the readers to [57] for more details about the tracking. This kind of conditions makes the robot face simulated occlusions and estimation errors.
@@ 4.2.4 Reward function
where $d_{t,goal}=||\mathbf{g}-(x_{t},y_{t})||$ and $\Delta{d_{t,goal}}=d_{t,goal}-d_{t-1,goal}$. The robot receives a reward of $r_{goal}=15$, when it reaches the goal within a certain threshold; and a negative reward $r_{collision}=-15$ when it collides. Reward shaping is used to accelerate training in non-terminal states by encouraging the agent to get closer to the goal ($-r_{dist}\Delta{d_{t,goal}}$ with $r_{dist}=2.5$), and by penalizing the agent if it is too close to an obstacle with:
@@
4.3 Network
We propose using a Soft Actor Critic (SAC) [23] algorithm to solve the problem, and use the DOVS model as part of the input of the network to extract the dynamism of the environment. The implementation of Stable Baselines 3 [60] is used for the SAC. SAC is an off-policy DRL algorithm that uses an actor to determine the policy and a critic to estimate the state-action values, and introduces entropy maximization to help exploration; it is known for its stability and sample efficiency.
@@ 4.3 Network
$\mathbf{o}_{t,state}=\left\{\hat{\mathbf{u}}_{t},\hat{d}_{t,goal},\hat{\phi}_{t,goal},\hat{\mathscr{O}}_{t,dist},\hat{\mathscr{O}}_{t,\theta},\hat{\mathscr{O}}_{v,t},\hat{\mathscr{O}}_{t,dir}\right\},$ (23)
where $\hat{d}_{t,goal}$ and $\hat{\phi}_{t,goal}$ are the distance and angle to the goal with respect to the estimated heading angle of the robot ($\theta_{t}$), $\hat{\mathscr{O}}_{t,dist}$ the estimated distance of the robot to the closest obstacle, $\hat{\mathscr{O}}_{t,\theta}$ the estimated angle to it, $\hat{\mathscr{O}}_{v,t}$ its estimated velocity and $\hat{\mathscr{O}}_{t,dir}$ its estimated heading angle. Even though the information of these four last variables is already encoded in DOVS, they are included to give more information to the robot in case there is an imminent collision.
@@
The output of both streams is concatenated and fed to an LSTM layer [43], $\psi_{LSTM}$, to save information of previous observation and keep temporal dependencies. The future of the environment is already considered by DOVS, but considering the past could help the network understand changes in the obstacles behavior and make it more robust to estimation errors or occlusions. Finally, a fully connected layer with ReLU activation function, $\psi_{FC}$, produces the final encoded observation $\mathbf{o}_{t,enc}$:
@@ 5 Experiments
This section details RUMOR evaluation, both with simulation experiments that compare it with existing motion planners (Section 5.2 and hardware experiments that prove its performance in a ground robot (Section 5.3.
5.1 Experimental setup
Simulation setup. The planners are tested in a modified version of the Stage simulator [61]. It is also used for training RUMOR and the DRL baselines. This simulator is conveniently integrated in ROS to use AMCL for localization and realistically simulates a LiDAR sensor. The scenario used is an 6x6 m open space with dynamic obstacles the robot has to avoid to reach the goal. The scenario has walls far from this space to let AMCL localize the robot. The obstacles try to avoid each other using the ORCA algorithm, but cannot see the robot to make it perform the whole obstacle avoidance maneuvers, as stated in Section 4.2.3. Random positions are sampled for the robot, its goal and the other obstacles. The obstacles have a predefined random linear and angular velocity that try to follow. The linear velocity of obstacles is sampled between $\frac{v_{max}}{5}$ and $v_{max}$ m/s, where $v_{max}$ is the robot’s maximum velocity. This range ensures variability in the obstacle behavior. The robot velocity limits are set to $v_{max}=0.7$ m/s and $\omega_{max}=\pi$ rad/s and the maximum acceleration to $a_{max}=0.3$ m/s², similar to those of a Turtlebot 2 platform. Although typical human walking speeds exceed the obstacles velocity range, setting it to match the robot’s maximum speed allows for effective training in obstacle avoidance, despite the TurtleBot 2 being relatively slow. Moreover, the speed of people walking in crowds and when facing other people in open spaces is usually lower than the typical values.
We conducted the simulation experiments with two different perception settings: using the ground truth position of the dynamic obstacles given by the simulator (absolute perception) and using the Kalman Filter-based obstacle detector [57] (obstacle tracker) previously introduced in Section 4.2.3, thus facing occlusions and estimation errors.
Training setup. During training, the robot faces scenarios with increasing number of dynamic obstacles, up to 14, and increasing starting distance to the goal during a first training stage of 1000 episodes. In this way, it learns to reach the goal progressively by using curriculum learning, from simple to more complex tasks. After that stage, it faces 9000 scenarios with a random number of obstacles up to 14. The minimum distance between the robot and the goal is set to 6 m.
@@ 5.1 Experimental setup
All DRL policies (both RUMOR and those used as baselines) have been trained with this setup. In addition, they have been trained in both perception settings to properly train them, i.e., there are two versions of each of the DRL planners, one trained using absolute perception and the other using the obstacle tracker.
Hardware setup. RUMOR was integrated in a Turtlebot 2 platform with a NUC with Intel Core i5-6260U CPU and 8 GB of RAM. The sensor used is a 180º Hokuyo 2D-LIDAR. Due to the real-world design of the system, the same network weights and ROS nodes used in simulation were used in the ground robot, with the same Kalman Filter-based obstacle detector and AMCL for localization. Therefore, the RUMOR version used in the simulation experiments with the obstacle tracker is the same one as the used as in the hardware ones, with the same network weights. Obstacles like walls are detected as segments by the obstacle detector. They are included in the DOVS as colliding velocities by simply computing the velocities that intersect with the segment within a time horizon. This allows the robot to simply detect them without modifying the design.
To help the robot navigate along long-paths, through rooms and avoid other non-convex obstacles, we use a way-point generator introduced in [63] to feed the robot with intermediate goals rather than simply sending the final goal to it. The idea behind it is first computing a global plan with A regarding the static map of the environment and filter the plan to get sparse intermediate points that are corners of sharp turns. A new goal from this set of intermediate goals is sequentially set as the RUMOR goal when the robot is close to approaching the previous one, letting RUMOR compute the motion commands by itself but with intermediate guidance that helps it to avoid convex obstacles.
@@ 5.2 Simulation experiments
We conducted quantitative experiments to evaluate the performance of our proposed model. We compared our method, RUMOR, with DWA [5], TEB [9], ORCA [15], SGAN-MPC [64], S-DOVS [1], SARL [45] , RGL [46], RE3-RL [48] and SG-D3QN [52].
Remark 2.
The comparison of RUMOR with the baselines planners is unfair, as they do not consider the same velocity constraints. ORCA, SGAN-MPC, SARL, RGL, RE3-RL and SG-D3QN only consider maximum velocity limits, overlooking constraints of Equations 3 and 4. For a more fair comparison, we consider a different version of RUMOR with no restrictions (NR-RUMOR). NR-RUMOR has a continuous action space that comprises the whole robot velocity range: $\mathbf{a}_{t}=(a_{1,t},a_{2,t})\in[0,v_{max}]\times[-\omega_{max},\omega_{max}]$.
Remark 3.
We used ROS navigation stack implementations of DWA and TEB, which considers acceleration restrictions in $v$ and $w$ separately. As shown before, the maximum linear acceleration/velocity of a differential-drive robot is directly related to its angular acceleration/velocity. RUMOR accounts for these restrictions, but DWA and TEB do not. The result of this difference may be understood with Figure 2. First, the restrictions implemented in DWA and TEB do not consider kinematic constraints of Equation 3, so their velocities may surpass the two black lines. Second, the acceleration constraints would be equivalent to, instead of using the rhombus of Equation 4 as the dynamic window around the current robot velocity, using a bounding box that surrounds the rhombus. DWA and TEB therefore have an acceleration and velocity space that doubles the size of RUMOR’s, which accounts for every differential-drive constraint.
Remark 4.
We selected as agent-based DRL baselines SARL, RGL, RE3-RL and SG-D3QN. Even though we may adapt RUMOR to their restrictions (NR-RUMOR), these baselines may not be adapted to consider the differential-drive constraints. First, they use a discrete action space, so RUMOR action space is not available for them. Second, it is not clear how to restrict a discrete action space avoiding issues in Remark 3. Third, modifying those methods to use a continuous action space would imply changing their algorithm, which could alter training time or stability.
We tested the planners in the same 500 random scenarios with low and high occupancy, with 6 and 12 dynamic obstacles, respectively. The 85% of the obstacles are set to be dynamic and the same set of scenarios is used for every method.
We mainly evaluate the planners using the two navigation metrics directly encoded in the reward function, which are the success rate (number of times the robot have reached the goal without collisions or timeouts) and the navigation time. There is a trade-off between both metrics. Trying to achieve a low navigation time leads to risky collision avoidance maneuvers, which leads to lower success rates. Collisions are sometimes unavoidable, particularly in dense environments or when estimation errors occur. In such cases, the robot may encounter trapping situations from which it cannot escape or face obstacles that move unpredictably toward it. Thus, achieving consistently high success rates in complex scenarios is unrealistic, so these results should be interpreted as a benchmark. In real-world settings, where obstacles typically cooperate in collision avoidance or at least do not actively move toward the robot, the success rate is likely to be higher.
The success rates obtained in the experiments are presented in Table 1. Scenarios in which all methods failed were excluded from the evaluation metrics. To assess statistical significance, we performed a $\chi^{2}$ test to compare the success rate of NR-RUMOR and RUMOR against the baselines. In addition, we recorded the time to reach the goal, mean velocity and path length metrics for each testing episode where all methods succeeded. The results are visualized as box-plots in Figure 7 for DWA and the DRL planners (the rest are not included to simplify visualization). Furthermore, we conducted Mann-Whitney U rank tests with a one-sided (”less”) alternative hypothesis to statistically determine if RUMOR and NR-RUMOR achieved significantly lower navigation times compared to the other methods.
Method Abs. Per. Obs. Track.
6 Obs. 12 Obs. 6 Obs. 12 Obs.
TEB 0.47∗† 0.10∗† 0.47∗† 0.10∗†
DWA 0.69∗† 0.44∗† 0.70 0.46
S-DOVS 0.70∗† 0.36∗† 0.58∗† 0.37∗†
SGAN-MPC 0.66∗† 0.51∗† 0.56∗† 0.36∗†
ORCA 0.66∗† 0.51∗† 0.60∗† 0.35∗†
SARL 0.81∗† 0.64∗† 0.74 0.46
RE3 0.87∗† 0.71 0.64∗† 0.35∗†
RGL 0.88 0.68 0.65 0.39∗†
SG-D3QN 0.91 0.78 0.72 0.50
RUMOR 0.91 0.72 0.70 0.46
NR-RUMOR 0.94 0.82 0.78 0.58
Table 1: Success rates for the different methods for the simulation experiments, using absolute perception and the obstacle tracker. The method with the best score in each setup is in bold. denotes statistical significant difference between NR-RUMOR and a method, and ∗† between both NR-RUMOR and RUMOR and a method, with a p-value $<0.05$.
Absolute perception (6) Absolute perception (12)
@@
(a1) Navigation time (a2) Navigation time
@@
(b1) Path length (b2) Path length
@@
(c1) Mean v (c2) Mean v
Obstacle tracker (6) Obstacle tracker (12)
@@
(a3) Navigation time (a4) Navigation time
@@
(b3) Path length (b4) Path length
@@
(c3) Mean v (c4) Mean v
@@
Figure 7: Navigation metrics for dynamic scenarios of 6 (left) and 12 (right) obstacles using absolute perception and the obstacle tracker. In (a), denotes statistical significant difference between NR-RUMOR and a method, and ∗† between both NR-RUMOR and RUMOR and a method, with p-value $<0.05$.
Overall, NR-RUMOR has the highest success rate in all settings, while having lower or at least comparable navigation times. Moreover, the difference with the rest of the planners is statistically significant in most of cases. The only case where the difference is not significant in both low and high density scenarios is when comparing NR-RUMOR with SG-D3QN in absolute perception settings. However, the mean time taken by NR-RUMOR planner to reach the goal is significantly lower in those cases.
@@
The success rates show that the agent-model based planners (S-DOVS, SGAN-MPC and ORCA) performance is worse than DRL-based ones regarding their success rates. These planners use the same agent information DRL planners as input. They model the environment in every time step, but they are not robust when errors or deep changes in the surroundings occur.
@@
Apart from TEB and DWA, which do not use the agent information as input, every planner shows higher success rates and lower navigation times with absolute perception, due to the absence of partial observability and estimation errors. It is way more complex to predict the future of the environment when facing perception problems. Therefore, their relative performance within the same settings should be compared. It is interesting to see that there is an inconsistency in the performance of SARL and RE3-RL or RGL when changing the perception settings. When comparing them using absolute perception, RE3-RL and RGL outperform SARL in both success rate and navigation time, while when using the obstacle tracker is the other way around. This is probably due to the difference in the complexity of their networks and the exploration process, degrading the performance of complex models in the presence of erroneous observations. In contrast, both NR-RUMOR show a very consistent relative performance no matter the perception settings. SG-D3QN shows a consistent performance too, but, as stated before, with weaker results than NR-RUMOR.